OpenStudy (anonymous):

Find all real zeros of the function y = 7x + 4 -4/7 -7,4 -7 4/7

OpenStudy (anonymous):

y = 7x + 4 Let y = 0 and solve for x

OpenStudy (anonymous):

B @TURITW

OpenStudy (anonymous):

It's A 7x + 4 = 0 7x = -4 x = -4 / 7

OpenStudy (anonymous):

thanks now I know how to do #2